OpenStudy (umulas):

1. Alice had to read 350 pages of a book over the weekend. If on Sunday, she read 50 pages more than half the amount she read on Saturday, how many pages did she read on SaturdaY? (A) 150 (B) 175 (C) 200 (D) 225 I got in my book C, but I would like to know WHY.

OpenStudy (anonymous):

What she read on Saturday:x 50+1/2 x=350

OpenStudy (anonymous):

That's funny

OpenStudy (anonymous):

I got the wrong answer never mind

OpenStudy (anonymous):

Here, there is another way to solve it Ok

OpenStudy (umulas):

I mean I got Sun: 50 + 1/2x Sat: x = 350 50+1 1/2 x = 350 For some people, I keep finding 3x1/2, I don't know why though.

OpenStudy (umulas):

For some reason: 50 + 3x1/2 = 350 3x1/2= 300 3x=600 x=20 I just don't know where they got 3 : P

OpenStudy (anonymous):